OpenStudy (anonymous):

RELATED RATES QUESTION! Gravel is being dumped from a conveyor belt at a rate of 30 cubic feet per minute. It forms a pile in the shape of a right circular cone whose base diameter and height are always the same. How fast is the height of the pile increasing when the pile is 12 feet high?

OpenStudy (isaiah.feynman):

Cone...what equation relates the volume, radius, and height?

OpenStudy (anonymous):

\[1/3*\pi*r^2*h\]

OpenStudy (isaiah.feynman):

You are a listening student. Now, differentiate it implicitly.

OpenStudy (anonymous):

can i replace the r with h/2 since the problem says that D and h are always the same?

OpenStudy (isaiah.feynman):

Try that.

OpenStudy (isaiah.feynman):

Yes! you are thinking, replace r with h/2 so we don't have two unknowns to solve for! I didn't even see that at first.

OpenStudy (isaiah.feynman):

Wait a minute let me solve.

OpenStudy (anonymous):

ok(:

OpenStudy (ranga):

I am getting dh/dt = 0.2653 feet / min. when h = 12 feet.

OpenStudy (anonymous):

thats right! how did you get that? :O

OpenStudy (ranga):

V = 1/3pir^2h r = h/2 V=1/3pi(h^2)/4 * h V=pi/12h^3 dV/dt = pi/12 * 3h^2 * dh/dt 30 = pi/4h^2 * dh/dt h = 12 solve for dh/dt

OpenStudy (anonymous):

\[V=\frac{ 1 }{ 3 }\left( \frac{ h }{ 2 } \right)^2h \rightarrow V=\frac{ \pi h^3 }{ 12 }\rightarrow \frac{ dV }{ dt }=\frac{ 3\pi h^2 }{ 12 }\frac{ dh }{ dt }\] Did you derive it correctly?

OpenStudy (anonymous):